OpenStudy (anonymous):

The length of a rectangle is 4 inches longer than its width. What are the possible widths if the area of the rectangle is at least 896 square inches?

OpenStudy (anonymous):

Anybody?

OpenStudy (anonymous):

lw=896 l=w+4 w(w+4)=896 w^2+4w=896 w=28 l=28+4 l=32

OpenStudy (anonymous):

I've answered it!

OpenStudy (anonymous):

32 is not the answer. I tried earlier.

OpenStudy (anonymous):

no, 28*32=896.

OpenStudy (anonymous):

ok thanks